Shift report 20000723 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Jul 23 2000 - 08:13:03 EDT

  • Next message: http account: "Shift report 20000723 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Jul 23 08:13:02 2000, Shift: 00:00-08:00
    Supervisor: FR, with: RP
    Log book#3 pages 48 - 52
    
    Shift Summary:
    ------------------------------------------------------------
    
    -There was a successful ramp at 2:30. Had a long store with a cogged 
     beam for about 2 hours.Intensity of the yellow was 
     relatively low: yellow = 1.3*10^8, blue = 8.3*10^8.
     Had another store at 5:20 but it lasted only for 20 minutes.
     Intensity of yellow beam was slightly higher: 1.6*10^8.
     At 7:25, we had another store for about 30 minutes 
     (intensity of yellow beam: about 2*10^8, blue: 8.4*10^8).
    
    -Three good runs at gamma=70.  
     Run1889 : about 110 minutes, contains about 33 collisions.
     Run1890 : about 20 minutes, 25 collisions.   
     Run1891 : about 30 minutes, 40 collisions.
     Data rate was very low (1-2/sec.).
    
    -HV control has still problems connecting to H1,H2,C1,TOFW.
     No other problems to report.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Jul 23 2000 - 08:16:17 EDT